OpenStudy (anonymous):

Find y' and y'' y=e^9e^x I'm lost and don't know what to do here.

OpenStudy (anonymous):

\(\large y=e^{\Large9}e^x\) or \(y=e^{\large9e^x}\) ?

OpenStudy (anonymous):

The second one. Sorry, I'm not familiar with the formatting.

OpenStudy (anonymous):

That's okay\[\begin{align} y' = e^{\Large9e^x}\cdot9e^{\large x} &= 9e^\left({\Large{9e^x+x}}\right) \\ y'' = (y')' &= 9e^\left({\Large{9e^x+x}}\right)\cdot(9e^{\large x}+1) \end{align}\]

OpenStudy (anonymous):

Thanks so much! I was lost to as to how to get it.